OpenStudy (anonymous):

Use points (1,-2) and (2,7) and put into point slope form.

OpenStudy (anonymous):

Point slope form is: \[y-y ^{1} = m(x-x ^{1})\]

OpenStudy (anonymous):

*And find the equation of the line

OpenStudy (anonymous):

Go ahead and plug your points in.

OpenStudy (anonymous):

So -2-7=m(1-2)

OpenStudy (anonymous):

Yes. That is point-slope form. However, we now need to find the equation.

OpenStudy (anonymous):

This can be done by simply graphing. I'll graph and send a link.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

https://www.desmos.com/calculator/y4l072lo1c Use this one

OpenStudy (anonymous):

The equation of this line can be put into y=mx+b To do so, we can use the points to find the slope (m) and the graph to find the y int (b)

OpenStudy (anonymous):

Follow?